# Edge Weight Prediction For Category-Agnostic Pose Estimation

### v1

@@ 4.4 Ablation Study
Following standard practice, we conducted several ablation studies on the MP-100 split-1 in the 1-shot setting. First, we present a quantitative comparison in the challenging super-category setting. Next, we analyze the contribution of each component within our framework. Finally, we examine the impact of different input graphs $A_{prior}$ on performance.
In the supplementary, we provide additional ablation experiments including masking query images to evaluate occlusion handling, quantitative analysis and comparisons of our adjacency matrix supervision strategy, histograms of predicted adjacency weight changes, and evaluations with various backbones.
@@
POMNet [42] 73.82 79.63 34.92 47.27
CapeFormer [35] 83.44 80.96 45.40 52.49
ESCAPE [24] 80.60 84.13 41.39 55.49
MetaPoint+ [7] 84.32 82.21 46.51 53.67
SDPNet [31] 83.84 81.24 45.53 53.08
PPM [27] 85.13 82.44 52.08 60.59
SCAPE [19] 84.24 85.98 45.61 54.13
GraphCape [14] 88.38 83.28 44.06 45.56
Ours 88.51 89.62 45.55 64.80
Table 2: Super Cross-Category. PCK results for the super cross-category setting. Our method outperforms others in this challenging setting across most splits, demonstrating robust generalization. The best results are bold, and the second-best are underlined.
@@
We conduct a cross-super-category experiment following prior works to assess our model’s generalization capacity. Splitting the MP-100 dataset into eight super-categories, we hold out each super-category in turn and train on the rest. As shown in Table 2, our model outperforms other methods across most splits. We attribute much of this improvement to our proposed feature extraction, which combines a frozen DinoV2 with a DPT decoder. This harnesses its fine-grained features for precise localization without impairing generalization. This outcome highlights our framework’s effectiveness in capturing category-agnostic features.
Method Ablation PCK0.2
GraphCape [14] 91.19
+ Dino Backbone 90.34 (-0.85%)
+ Dino Probing 92.28 (+1.94%)
+ Remove Heatmap Loss 92.80 (+0.52%)
+ Graph Weight Prediction 93.54 (+0.74%)
+ Attention Bias 93.69 (+0.15%)
Table 3: Method Ablation. We showcase the contribution of each of our modified components. Simple backbone replacement reduces accuracy, while learnable multi-resolution fusion, attention bias, and edge weight prediction each boost performance, highlighting the value of complex weighted graphs for localization.
Components Contribution.
We base our architecture on GraphCape [14], with Table 3 demonstrating the performance impact of each proposed component as they are added incrementally. Replacing the backbone initially results in a drop in accuracy, but incorporating our suggested feature extraction method yields a substantial performance gain. Noticeable improvements are achieved by integrating our graph prediction model, with an additional boost provided by the Markov Attention Bias. Together, these results underscore the value of using complex weighted graphs for accurate localization.
Prior Graph Input $A_{prior}$ PCK0.2
Self-Connections Only 93.10
Fully-Connected Graph 92.91
Annotated Graph 93.69
Table 4: Different Prior Graph Inputs. Results using different $A_{prior}$ inputs. Testing with disconnected or fully connected graphs shows a performance drop, as the model is designed to refine rather than fully predict the structure.
The Impact of Prior Graph Input $A_{prior}$.
To evaluate the impact of the input graph structure prior, we test our method, trained on annotated graphs, using different graph inputs. Results are shown in Table 4. Evaluating the model with disconnected or fully connected graphs leads to a performance drop, as our model is designed to refine the structure rather than predict it entirely. In the supplementary material, we provide qualitative examples and illustrate changes in adjacency matrix weights for different $A_{prior}$.
5 Conclusion
We present EdgeCape, a novel CAPE method that improves keypoint localization through predicted refined skeleton graphs. Unlike previous methods that treat keypoints as isolated or use unweighted graphs, EdgeCape predicts weighted graphs that yield superior localization accuracy.
Our approach integrates edge weight prediction with Markovian Attention Bias to capture complex structural dependencies, improving accuracy under occlusions and symmetry challenges. Evaluated on the MP-100 dataset, EdgeCape achieves state-of-the-art results in the 1-shot setting and leads among similar-sized methods in the 5-shot setting, advancing versatile, structure-aware CAPE methods.
Supplementary Material
Appendix A Supplementary Index
  • Additional metrics results on MP-100 dataset.
  • Quantitative evaluations on a single-category setting.
  • Method scalability - performance using different backbone models in various sizes.
  • Query image masking performance comparison, demonstrating occlusion handling.
  • Adjacency matrix supervision strategy ablation study.
  • Graph prediction module design ablation - the importance of support image features for graph prediction.
  • More qualitative results - including predicted graphs, and qualitative comparisons.
  • Implementation Details.
  • Framework Overview.
  • Markovian Attention Bias - Complete description and derivation.
  • Training Scheme.
Appendix B Further Experiments
B.1 Quantitative Results - Additional Metrics
Model AUC $\uparrow$ NME $\downarrow$
CapeFormer [36] 88.64 86.39 86.18 85.81 86.51 86.70 0.088 0.110 0.111 0.116 0.108 0.106
SCAPE (DinoV2) [19] 89.85 87.41 87.85 87.09 87.72 87.98 0.074 0.099 0.093 0.096 0.096 0.092
GraphCape [14] 89.08 87.69 86.97 87.01 86.67 87.48 0.083 0.097 0.103 0.104 0.107 0.099
Ours 90.88 88.35 88.17 87.69 88.00 88.62 0.064 0.090 0.090 0.097 0.093 0.088
For a more complete comparison of our method, we report two additional metrics - Area Under ROC Curve (AUC) and Normalized Mean Error (NME) on the MP-100 dataset 1-shot setting. We evaluate CapeFormer, SCAPE, GraphCape, and our method. As shown in Table 5, our method achieves state-of-the-art also in these metrics, demonstrating the superiority of our method.
B.2 Ablation Study
Model Metric
PCK  $\uparrow$ AUC  $\uparrow$ NME  $\downarrow$
CapeFormer [36] 87.44 86.41 0.094
GraphCape [14] 89.54 87.22 0.086
Ours (SwinV2) 89.90 87.29 0.085
Ours (DinoV2 + DPT) 89.93 87.63 0.081
Table 6: AP-10K Results. PCK, AUC, and NME results for the single-category AP-10K benchmark. [] methods fine-tune their backbone. The best results are bold.
Single-Category Setting.
We report PCK performance on a single-category pose estimation benchmark using the AP-10K dataset [49], a large-scale animal pose estimation benchmark comprising 23 animal families and 60 species. This dataset features a unified keypoint definition and skeletal structure across all categories. Notably, the training and testing category splits are not mutually exclusive, meaning categories may appear in both sets. This setup allows us to evaluate our method’s effectiveness in a single-category context, where categories share common anatomical structures but exhibit species-specific variations.
The results are presented in Table 6. Our method outperforms CapeFormer [36] and GraphCape [14]. Notably, the compared methods fine-tune their backbones, which provides an advantage in this single-category setting where only a fixed set of keypoints is relevant. We present two versions of our method: one using a fine-tuned SwinV2 backbone (as in GraphCape) and another employing the feature extractor proposed in this paper. As the results show, leveraging predicted weighted graphs and Markovian Attention Bias leads to a performance gain. Moreover, although our proposed feature extractor uses a frozen backbone, it achieves slightly better performance than the SwinV2 fine-tuned version.
@@ B.2 Ablation Study
We assess how the weights of the adjacency matrices change. We create a histogram of the difference between normalized $A_{prior}$ and the normalized predicted $A$ matrix. Results are shown in Figure 8, demonstrating that our network effectively refines the input graph prior. This refinement involves creating new connections and adjusting the weights of existing ones, enabling the model to better represent structural relationships.
Backbone GraphCape [14] Ours
SwinV2 91.19 91.52
DinoV2 + DPT 92.80 93.69
DinoV2-B + DPT - 94.10
Table 7: Different Backbones. PCK results using different backbones for GraphCape [14] and our method. Our method outperforms GraphCape with the same backbone and demonstrates scalability with larger backbones. Results also highlight the benefits of Dino features, which exhibit 3D structural awareness.
Method Scalability.
We assess our model’s performance using backbones of varying sizes to explore the trade-off between accuracy and efficiency. Results are shown in table 7. As anticipated, larger backbones yield improved performance due to their increased capacity for feature extraction and representation. However, these improvements come at the cost of greater computational requirements and model size. For comparison, we include results using GraphCape’s [14] original SwinV2 backbone. This comparison underscores the advantages of our weighted graph refinement approach. Nevertheless, as hypothesized in the main paper, the performance gap is less pronounced because Dino features, utilized in our model, offer inherent 3D structural awareness. This capability is crucial for accurately predicting and refining structural relationships, making Dino a particularly suitable choice for our task.
Figure 9: Handling Occlusions. Quantitative results when masking the query image. Our method consistently surpasses GraphCape, leveraging cues provided by the weighted graph structure to overcome information gaps in the query images.
Handling Occlusions.
To demonstrate the effect of weighted graph information in handling occlusions, we applied random partial masking to the query images before executing our algorithm. As shown in Figure 9, our method consistently surpasses GraphCape when parts of the query images are masked, accurately predicting keypoints.
Adjacency Matrix Supervision.
To evaluate the effectiveness of our unsupervised masking strategy, we first test the performance using the GCN reconstruction approach proposed by Ren et al. [31], which utilizes an auxiliary GCN to reconstruct masked inputs. With this approach, we achieve 92.89% accuracy, compared to 93.69% using our decoder-based reconstruction strategy — a decrease of 0.8%. This drop supports our hypothesis that structural information is embedded within the auxiliary GCN weights and thus lost during inference.
@@
Graph Prediction Module Design.
We highlight the importance of utilizing whole-image features for skeleton prediction rather than just keypoint features. Incorporating cross-attention with support image features in the skeleton prediction network achieves 93.54% PCK. In contrast, excluding image features results in a performance drop to 93.43% (-0.11%), demonstrating the value of leveraging global context for accurate graph refinement.
B.3 Additonal Qualitative Results
@@
Support GT CapeFromer [36] SCAPE [19] GraphCape [14] EdgeCape
@@
Figure 10 illustrates the instance adaptability of our graph prediction module. Specifically, we present two input graphs of a panda body: one where all keypoints are visible, and another where some keypoints are occluded, resulting in isolated nodes in the input graph. As shown, our graph prediction module generates different graphs for instances of the same category based on the input. Notably, the predicted graph connects the isolated node, enhancing localization accuracy. In addition, Figure 11 illustrates additional skeleton predictions. We also present in Figure 12 additional qualitative comparison of Capeformer [35], SCAPE [19], GraphCape [14], and our model on various categories.
B.4 Implementation Details
For a fair comparison, training parameters, data augmentations, and data pre-processing are kept the same as in previous works. In addition, the backbone size was matched to GraphCape, thus we used the smallest (20M parameters) version of DinoV2. The graph predicting network, encoder, and keypoint prediction decoder have 3 layers. For the Markovian Bias Attention, we use a maximum of $K=4$ hops, and for $\mathcal{L}_{adj}$ we mask 50% of keypoints, using $\lambda_{adj}=1$. The model is built upon MMPose framework [8], trained using Adam optimizer with a batch size of 16, the learning rate is $10^{-5}$, and decays by 10× on the 160th and 180th epoch. Each phase is trained for 100 epochs. Training takes $\sim$ 1 day on a single Nvidia A100 GPU. The feature extraction DPT decoder is based on the official code from [10]. Evaluation of SCAPE [19] was done by removing the keypoint identifiers (which were shown to be inapplicable for real-world scenarios [24, 44, 14]) and training a network using their official code. We also used GraphCape and CapeFormer official code for evaluation and visualizations.

### v2

@@ 4.3 Ablation Study
Table 3: Method Ablation. We show mPCK results demonstrating the contribution of our category-agnostic pose-graph prediction module and the Markov Attention Bias.
Markov
Attention Bias
Pose-Graph Prediction (-) 79.87 80.78
(+) 80.76 81.96
@@
Following standard practice, ablation studies are conducted on MP-100’s split-1 in the 1-shot setting. We analyze the contribution of each component within our framework and examine the impact of input pose-graphs $A_{prior}$ on performance. In the supplementary (Section B.2), we provide additional extensive ablation experiments, including a quantitative comparison in the challenging super cross-category setting, performance using different backbones, evaluation without $A_{prior}$, evaluation of occlusion handling, quantitative analysis of different design choices and our suggested supervision strategy, and histograms of predicted adjacency weight changes.
Components Contribution.
Table 3 shows the importance of both the Markov Attention Bias and the category-agnostic pose-graph prediction model. We build on GraphCape (Hirschorn and Avidan, 2024), incorporating each component separately. The results show that both components contribute to improved performance, with the greatest gain achieved when they are combined. The Markov Attention Bias not only improves performance on its own but also benefits further from the enhanced predicted pose-graphs, resulting in an even greater overall boost.
Figure 6: mPCK with noisy $A_{\text{prior}}$. Testing on graphs with randomly added edges shows that our method is robust to incorrect graph inputs, effectively predicting pose-graphs for improved localization.
The Impact of Prior Graph Input $A_{prior}$.
To evaluate the impact of the input pose-graph prior, we compare our method to GraphCape using noisy graph inputs $A_{prior}$, as shown in Figure 6. We inject up to 16 random edges per pose-graph, comparable to the original skeleton’s edge count, spanning noise levels from mild to fully randomized topologies. Since graph-based models rely on connectivity for localization, incorrect connections mislead the model by propagating cues between unrelated keypoints, degrading performance. Yet, our method consistently outperforms GraphCape, with the gap widening as more random edges are added. This highlights our model’s resilience to input pose-graph variations and its ability to predict optimal pose-graphs. Moreover, since the Markov Attention Bias heavily depends on the pose-graph, our superior performance further underscores the effectiveness of our prediction module. This robustness is especially valuable in real-world scenarios where defining an optimal pose-graph is difficult.
5 Conclusions
We introduce EdgeCape, a CAPE method that improves keypoint localization via predicted category-agnostic pose-graphs. Unlike previous approaches that treat keypoints independently or rely on unweighted user-provided graphs, EdgeCape predicts weighted pose-graphs for more accurate localization. By integrating these predictions with Markov Attention Bias, our model captures complex structural dependencies, improving robustness to occlusions and symmetry. On MP-100, EdgeCape achieves state-of-the-art performance in 1-shot and 5-shot settings. Beyond CAPE, our framework highlights the broader potential of weighted graph refinement for category-agnostic vision tasks under data scarcity and structural variability.
Appendix A Supplementary - Introduction
  • Additional MP-100 quantitative evaluations.
  • Quantitative evaluations on the AP-10K dataset.
  • Super-cross category experiment, emphasizing generalization.
  • Occlusion handling experiment, using query image masking performance comparison.
  • Supervision strategy ablation studies.
  • Additional qualitative results - including predicted graphs, and qualitative comparisons.
  • Framework Overview.
  • Markov Attention Bias.
  • Training Scheme.
  • Implementation Details.
Appendix B Further Experiments
Table 4: MP-100 PCK@0.1 Results. comparison between graph-based methods on the MP-100 dataset. Our method consistently outperforms others across all settings and data splits.
Method 1-Shot 5-Shot
CapeFormer 74.48 68.39 67.99 68.18 68.86 69.58 81.46 77.12 76.28 76.71 71.86 76.69
GraphCape 76.17 70.05 71.15 68.89 68.95 71.04 81.99 77.22 75.26 75.99 76.28 77.35
Ours 79.29 73.82 73.1 71.16 72.19 73.91 83.32 78.42 76.29 76.82 77.29 78.43
Model AUC $\uparrow$ NME $\downarrow$
CapeFormer 88.64 86.39 86.18 85.81 86.51 86.70 0.088 0.110 0.111 0.116 0.108 0.106
GraphCape 89.08 87.69 86.97 87.01 86.67 87.48 0.083 0.097 0.103 0.104 0.107 0.099
Ours 90.05 87.95 88.43 87.18 87.40 88.20 0.074 0.094 0.089 0.103 0.103 0.093
B.1 Quantitative Results - Additional Metrics
MP-100
For a more complete comparison of our method, we report PCK at threshold 0.1 in Table 4. As shown in Table 4, our method outperforms GraphCape, former SOTA graph-based method by +2.87% on the 1-shot setting and by +1.08% on the 5-shot setting. Overall, we achieve gains of +4.33% and +1.74% over CapeFormer, a method that don’t use any prior graph structure, emphasizing the importance of this added refined prior. Notably, even in splits where GraphCape underperforms CapeFormer, our method comes on top, underscoring the advantage of graph-based methods over point-based ones when optimal pose-graphs are learned and effectively utilized.
Additionaly we include the Area Under ROC Curve (AUC), and Normalized Mean Error (NME) on the MP-100 dataset 1-shot setting. As shown in Table 5, our method also excels in these metrics, demonstrating the superiority of our method.
Figure 7: MP-100 Comparison. Average PCK@0.2 comparison with recent methods. The circle size indicates the model’s size. The x-axis and y-axis represent the 1-shot and 5-shot PCK, respectively (best is top right). Our approach leads in both settings.
Point-based methods
We also include a visual quantitative comparison of the MP-100 dataset with recent CAPE methods in Figure 7. It’s important to note that most other non graph-based CAPE works are complementary and parallel to ours, meaning that combining different methods could help build even stronger CAPE models.
Computational Complexity.
EdgeCape introduces a lightweight computational overhead, adding approximately 2.5M parameters and 2 GFLOPs to the GrapCape baseline (totaling $\approx$ 30 GFLOPs). By leveraging recent transformer optimizations on modern hardware, this translates to a modest inference latency increase of only $\approx$ 2 ms per image on an NVIDIA A5000. Notably, the matrix-power operations required for attention bias are executed using PyTorch’s GPU-optimized linear algebra kernels, incurring negligible cost (measured at $\approx$ 0.02 GFLOPs for 100 keypoints and $K=4$). While we acknowledge that current CAPE methodologies have not yet reached real-time frame rates, we plan to investigate further efficiency improvements in future work.
B.2 Ablation Study
Table 6: Backbones Ablation. Comparison between graph-based methods on the MP-100 dataset (1-shot, split 1) using different backbones and metrics. Predicted weighted graphs improve performance, increasing the gap in performance using stronger pre-trained features.
Method Backbone Metric
PCK@0.1 mPCK
CapeFormer SwinV2 74.05 77.77
GraphCape 76.44 79.80
Ours 77.66 80.63
CapeFormer DinoV2 74.48 78.44
GraphCape 76.17 79.87
Ours 79.29 81.96
Performance Using Different Backbones.
Table 6 shows results on the MP-100 dataset under 1-shot setting. As can be seen, we consistently outperform CapeFormer (Shi et al., 2023b) and GraphCape (Hirschorn and Avidan, 2024), highlighting the effectiveness of our approach. Moreover, contrary to GraphCape’s claim (Hirschorn and Avidan, 2024), DinoV2 features outperform SwinV2 in graph-based CAPE methods. While previous works (Lin et al., 2023; Shi et al., 2023b; Liang et al., 2024) have found other backbones more suitable for localization tasks, our results suggest that DinoV2’s superior generalization and semantic understanding make it especially well-suited for category-agnostic settings. Importantly, we achieve these results while keeping the DinoV2 backbone frozen, preserving its rich self-supervised representations rather than fine-tuning, as done in prior methods. Furthermore, the performance gap increases when predicting pose-graphs using DinoV2 features, aligning with findings by Banani et al. (El Banani et al., 2024), who demonstrated the intrinsic 3D structural awareness of DinoV2. This structural understanding is particularly beneficial for capturing pose-graph structures.
Table 7: AP-10K Results. PCK@0.1, AUC, and NME results for the single-category AP-10K benchmark.
Model Metric
PCK@0.1  $\uparrow$ AUC  $\uparrow$ NME  $\downarrow$
CapeFormer 69.70 86.41 0.094
GraphCape 72.01 87.22 0.086
Ours 72.48 87.29 0.085
Single-Category Setting.
We report PCK performance on a single-category pose estimation benchmark using the AP-10K dataset (Yu et al., 2021), a large-scale animal pose estimation benchmark comprising 23 animal families and 60 species. This dataset features a unified keypoint definition and skeletal structure across all categories. Notably, the training and testing category splits are not mutually exclusive, meaning some categories may appear in both sets. This setup allows us to evaluate our method’s effectiveness in a single-category context, where categories share common anatomical structures but exhibit species-specific variations.
Results are presented in Table 7. Our method outperforms CapeFormer (Shi et al., 2023b) and GraphCape (Hirschorn and Avidan, 2024), showing superiority also in this setting.
Table 8: Super Cross-Category. PCK@0.2 results for the super cross-category setting. Our method outperforms others in this challenging setting across most splits, demonstrating robust generalization.
@@ B.2 Ablation Study
POMNet (Xu et al., 2022a) 73.82 79.63 34.92 47.27
ESCAPE (Nguyen et al., 2024) 80.60 84.13 41.39 55.49
MetaPoint+ (Chen et al., 2024) 84.32 82.21 46.51 53.67
SDPNet (Ren et al., 2024) 83.84 81.24 45.53 53.08
SCAPE (Liang et al., 2024) 84.24 85.98 45.61 54.13
CapeFormer (Shi et al., 2023a) 83.44 80.96 45.40 52.49
GraphCape (Hirschorn and Avidan, 2024) 88.38 83.28 44.06 45.56
Ours 91.38 87.61 44.12 54.03
@@
We conduct a cross-super-category experiment following prior works to evaluate our model’s generalization capacity. Splitting the MP-100 dataset into eight super-categories, we hold out each super-category in turn and train on the remaining categories. As shown in Table 8, our model outperforms other graph-based methods across most splits. Notably, in the furniture split, GraphCape (Hirschorn and Avidan, 2024) performs worse than CapeFormer (Shi et al., 2023b), indicating that in this case, the input pose-graph impedes performance. However, our graph-prediction network not only surpasses GraphCape but also outperforms CapeFormer. This demonstrates that our model successfully learns pose-graphs that improve localization, even when the input pose-graphs are suboptimal.
Performance Without Prior Knowledge
To assess the model’s dependency on informative structural priors, we conducted an ablation test using fully-connected graphs, which essentially removes topology-based information. Under this setting GraphCape suffer a performance collapse, reverting to the baseline CapeFormer level Hirschorn and Avidan (2024). In contrast, EdgeCape demonstrates significant robustness, achieving a +1.29% mPCK improvement over CapeFormer even with uninformative fully-connected graphs. Notably, this performance is on par with what GraphCape achieves using its standard input prior structure. This confirms that our method provides distinct benefits even when the graph prior carries no useful information.
Adjacency Matrix Supervision.
To evaluate the effectiveness of our unsupervised masking strategy, we first test the performance using the GCN reconstruction approach proposed by Ren et al.(Ren et al., 2024), which utilizes an auxiliary GCN to reconstruct masked inputs. With this approach, we achieve (PCK@0.2) 92.88% accuracy, compared to 93.57% using our decoder-based reconstruction strategy - a decrease of 0.69%. This drop supports our hypothesis that structural information is embedded within the auxiliary GCN weights and thus not retained during inference.
Furthermore, we investigate the impact of our decoder reconstruction strategy under varying levels of input keypoint masking. The results of this analysis are illustrated in Figure 8. As shown, masking keypoints enhances the supervision signal for the adjacency matrix, leading to more robust and structurally meaningful graph representations.
Figure 8: Keypoint Masking Supervision. Effect of the proposed unsupervised masking strategy with varying percentages of masked input keypoints
Pose-Graph Predictor Design.
We also ablate different design choices for the pose-graph predictor. First, we check the contribution of using full-image features compared to just the keypoint features to predict pose graphs. As hypothesized, using only keypoint features results in a decrease of 0.24% compared to using all features. This validates that the full-image features’ global context helps infer structural relations more effectively.
Moreover, we ablate the strategy of predicting pose graphs using the refined features. We compare the following strategies:
  • Using the suggested cosine similarity:
    $\Delta A_{ij}=<F_{refined}^{i},F_{refined}^{j}>$ (11)
  • Using an MLP-based prediction:
    $\Delta A_{ij}=\mathrm{MLP}\left([F_{\text{refined}}^{i}\,\|\,F_{\text{refined}}^{j}]\right)$ (12)
  • Using multi-headed attention-based prediction:
    $\Delta A_{ij}=\mathrm{Linear}\left(\left[Q_{1}^{i}\cdot K_{1}^{j}\,\|\,\dots\,\|\,Q_{H}^{i}\cdot K_{H}^{j}\right]\right)$ (13)
All methods achieve similar performance. Thus, we opt for the simplest design and use cosine similarity as our preferred strategy.
Figure 9: Handling Occlusions. PCK@0.2 results when masking the query image. Our method consistently surpasses GraphCape, leveraging cues provided by the weighted graph structure to overcome information gaps in the query images.
Handling Occlusions.
To demonstrate the effect of weighted graph information in handling occlusions, we applied random partial masking to the query images before executing our algorithm. As shown in Figure 9, our method consistently surpasses others when parts of the query images are masked, accurately predicting keypoints.
@@
We assess how the weights of the adjacency matrices change. We create a histogram of the difference between normalized $A_{prior}$ and the normalized predicted $A$ matrix. Results are shown in Figure 10. demonstrate that our network effectively refines the input graph prior. This refinement involves both creating new connections and adjusting the weights of existing ones, enabling the model to better represent structural relationships.
B.3 Additional Qualitative Results
@@
Input Graph Predicted Graph Input Graph Predicted Graph
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Predicted Graphs with Various $A_{prior}$. We visualize the unnormalized graph outputs with various $A_{prior}$ inputs. The left column denotes the input $A_{prior}$ and the right column is the refined adjacency matrix. Line width corresponds to edge weight.
@@
Support GT CapeFromer SCAPE GraphCape EdgeCape
@@
Figure 11 shows the instance adaptability of our graph prediction module. As can be seen, our graph prediction predicts different graphs for instances of the same category given different inputs. Moreover, Figure 12 shows skeleton predictions for different input values of $A_{prior}$. We use self-loops, visualizing performance without any prior knowledge, and with random omission graphs, where we drop some of the edge connections randomly. In addition, Figure 13 illustrates additional skeleton predictions.
We also present in Figure 14 additional qualitative comparison for Capeformer (Shi et al., 2023a), GraphCape (Hirschorn and Avidan, 2024), and the non graph-based SCAPE (Liang et al., 2024) on various categories.